OpenStudy (anonymous):

find dy/dx given that xy^2-x^2y+x^3=4

OpenStudy (tommo_lcfc):

This is an example of implicit differentiation. You need to use the product rule to differentiate the first two terms and then add a dy/dx to the terms you are differentiating with respect to y. Differentiating gives \[2xy dy/dx + y^2 - 2xy + x^2 dy/dx + 3x^2 = 0\] If we put all the terms with dy/dx to the other side, we get \[2xy - y^2 = (2xy + x^2) dy/dx\] Therefore \[dy/dx = (2xy - y^2) / (2xy + x^2)\] There is a common factor in both terms so this can be re-written as \[dy/dx = y(2x+y)/x(2y+x)\]

OpenStudy (tommo_lcfc):

Sorry, sign error, the + x^2 should be a -x^2 so we get dy/dx = y(2x-y)/x(2y-x)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

xy^2 - x^2 y + x^3 = 4 ( 2xy *y +y^2) - (x^2 *y' + 2xy) + 3x^2 = 0 3x^2 + 2xy + y^2 + 2xy *y' + x^2 y' = 0 (x^2 + 2xy) y' = - ( 3x^2 + 2xy + y^2 ) => y' = - ( 3x^2 + 2xy + y^2 ) / (x^2 + 2xy)